If af(cotx)+bf(cotx)=x then find f'(cotx) 
1)1/a-b
2)sin^2x/a+b
3)sin^2x/a-b
4)sin*^2x/b-a

Grade:11

1 Answers

Aditya Gupta
5 years ago
i think there is a slight error in your ques in sign of b. still, i will solve it assuming it is correct
we see that f(cotx)= x/a+b
now, differentiate both sides wrt x and use chain rule
f’(cotx)*(-cosec^2x)= 1/a+b
so f’(cotx)= – sin^2x/(a+b)